# A framework for LISA population inference

### v2

@@ 1 Introduction
The hierarchical Bayesian formalism [1, 2, 3] is now routinely applied in gravitational-wave (GW) astronomy to deconvolve measurement errors and selection effects in order to measure the global properties of source populations. It is used extensively in the analysis of LIGO/Virgo/KAGRA (LVK) [4, 5, 6, 7] observations to infer the astrophysical properties of stellar-mass compact objects [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21], to measure cosmological parameters [22, 23, 24], and to perform tests of General Relativity [25, 26, 27, 28]. A similar framework is also applied to pulsar-timing-array (PTA) [29, 30, 31] data in order to infer the properties of supermassive black holes in the late Universe [32, 33, 34]. Future detectors, such as the Laser Interferometer Space Antenna (LISA) [35, 36], will require a new population inference paradigm, however. Crucially, while current LVK observations consist exclusively of individually resolved mergers of stellar-mass compact objects, and PTAs observe a stochastic background generated by the superposition of inspiralling supermassive black hole binaries, neither of these regimes will apply to LISA, which will instead observe a superposition of resolved and unresolved signals from the same astrophysical population.
@@
The hierarchical Bayesian formulations for population inference usually treat either (i) a set of statistically independent, individually resolved events, or (ii) a stochastic signal that can be related to the underlying source population. However, they do not accommodate situations where a single population contributes simultaneously to a resolvable population and to an unresolved confusion background - precisely the case for LISA GBs, and potentially for EMRIs. A formalism that includes both resolved events and stochastic backgrounds was proposed in the LVK context [73], but it does not treat resolved and unresolved sources in a fully consistent way. As a consequence, it does not account for the correlations that naturally emerge in a LISA-like Global Fit, particularly the fact that the number of resolvable events is not fixed and that resolving more or fewer sources directly impacts the inferred background. As LISA data-analysis tools are being developed, it is becoming essential to establish a unified formalism capable of treating such populations, both to guide methodological development and to define the data products that should be provided by the collaboration. This issue is also relevant for third-generation ground-based detectors, which will come online in the next decade, such as the Einstein Telescope [74] and Cosmic Explorer [75], which are expected to detect an astrophysical stochastic background [76, 77, 78, 79] in addition to individual events, and for pulsar timing arrays, which are expected to resolve individual sources in the near future [80, 81, 82, 83], particularly once the Square Kilometre Array [84] becomes operational.
In this paper, we develop a more general hierarchical Bayesian framework capable of processing the output of a Global Fit in order to describe populations that generate both a stochastic background and individually resolvable signals. We discuss the conditions under which our formulation reduces to the standard hierarchical Bayesian approach and how it differs from that of [73]. Finally, using a toy model Global Fit, we illustrate how the framework can be applied in practice and highlight the interplay between the stochastic and resolved components in constraining the underlying astrophysical population.
@@ 2 Current hierarchical Bayesian formalism
In the above, the second term dependence on the full data is only an approximation, as it includes also all of the resolved-source containing segments, $\{d_{c}\}_{n_{\rm det}}$. This is acceptable for current ground-based detectors because data chunks containing detectable events are much rarer than those without. We can therefore assume that these rare events do not affect the reconstruction of the stochastic noise, and conversely, that the noise can be considered known when these events are observed, and does not need to be constrained jointly. However, as we will show in App. A, this expression does not fully incorporate the information of resolved and unresolved sources self-consistently. It is a good approximation for the current situation, where we can only place upper bounds on the background.
@@ 3 General hierarchical Bayesian formalism
$p(d,\{\theta\}_{n},n,\Sigma|\Lambda)\propto p(d|\{\theta\}_{n},\Sigma)p(\Sigma|\{\theta\}_{n})\prod_{i=1}^{n}N(\theta_{i}|\Lambda)e^{-N}.$ (3.1)
The above equation states that the likelihood for the data $d$ depends on the events and on the noise properties, and that the noise properties can also depend on the events drawn from $N(\theta|\Lambda)$. We will ultimately marginalise over $\{\theta\}_{n}$, $n$ and $\Sigma$ to obtain the population likelihood. The noise parameters $\Sigma$ can also contain information about quantities that do not depend on the source population, e.g., instrumental noise, but we keep the same notation for generality.
@@ 3.1 Separable case
$p(d|\Lambda)\propto\sum_{n_{1}=0}^{+\infty}\sum_{n_{2}=0}^{+\infty}\frac{N^{n_{1}}N^{n_{2}}}{n_{1}!n_{2}!}e^{-N}\int{\rm d}\{\theta\}^{1}_{n_{1}}{\rm d}\{\theta\}^{2}_{n_{2}}{\rm d}\Sigma\ p(d|\{\theta\}^{1}_{n_{1}},\Sigma)p(\Sigma|\{\theta\}^{2}_{n_{2}})\prod_{i=1}^{n_{1}}p(\theta^{1}_{i}|\Lambda)\prod_{j=1}^{n_{2}}p(\theta^{2}_{j}|\Lambda).$ (3.2)
Defining
@@
$p(d|\Lambda)\propto\sum_{n_{1}=0}^{+\infty}\sum_{n_{2}=0}^{+\infty}\frac{N_{1}^{n_{1}}N_{2}^{n_{2}}}{n_{1}!n_{2}!}e^{-N_{1}-N_{2}}\int{\rm d}\{\theta\}^{1}_{n_{1}}{\rm d}\{\theta\}^{2}_{n_{2}}{\rm d}\Sigma\ p(d|\{\theta\}^{1}_{n_{1}},\Sigma)p(\Sigma|\{\theta\}^{2}_{n_{2}})\prod_{i=1}^{n_{1}}p_{1}(\theta^{1}_{i}|\Lambda)\prod_{j=1}^{n_{2}}p_{2}(\theta^{2}_{j}|\Lambda).$ (3.6)
@@ 3.2 General case
$p(d|\Lambda)=$
$\sum_{n_{1}=0}^{+\infty}\sum_{n_{2}=0}^{+\infty}\frac{N^{n_{1}}N^{n_{2}}}{n_{1}!n_{2}!}e^{-R}\int{\rm d}{\theta}^{1}_{n_{1}}{\rm d}{\theta}^{2}_{n_{2}}{\rm d}\Sigma\ p(d|{\theta}^{1}_{n_{1}},\Sigma)p(\Sigma|{\theta}^{2}_{n_{2}})\prod_{i=1}^{n_{1}}p(\theta^{1}_{i}|\Lambda)\alpha(\theta^{1}_{i},\Lambda)\prod_{j=1}^{n_{2}}p(\theta^{2}_{j}|\Lambda)(1-\alpha(\theta^{2}_{j},\Lambda)).$ (3.11)
@@
$p(\Sigma|n_{2},\Lambda)=\int{\rm d}\{\theta\}^{2}_{n_{2}}\ p(\Sigma|\{\theta\}^{2}_{n_{2}})\prod_{j=1}^{n_{2}}p_{2}(\theta^{2}_{j}|\Lambda),$ (3.14)
@@
$p(d|\Lambda)\propto\sum_{n_{1}=0}^{+\infty}\int{\rm d}\{\theta\}^{1}_{n_{1}}{\rm d}\Sigma\ p(d|\{\theta\}^{1}_{n_{1}},\Sigma)p(\Sigma|\Lambda)\prod_{i=1}^{n_{1}}p_{1}(\theta^{1}_{i}|\Lambda)\frac{N_{1}^{n_{1}}}{n_{1}!}e^{-N_{1}}.$ (3.16)
This is main result of this paper, the general equation to be used for hierarchical Bayesian inference for LISA data. We will illustrate this with a single population chosen to represent GBs, but this can be readily extended to multiple populations, for example massive black hole binaries and EMRIs. In practice, when running a Global Fit with RJMCMC, we have samples from $p(\{\theta\}^{1}_{n_{1}},n_{1},\Sigma|d)$. Using Bayes’ theorem, and denoting by $\pi_{\rm PE}(\{\theta\}^{1}_{n_{1}},n_{1},\Sigma)$ the prior used to obtain the posterior during initial sampling, we have
$p(d|\Lambda)\propto\sum_{n_{1}=0}^{+\infty}\int{\rm d}\{\theta\}^{1}_{n_{1}}{\rm d}\Sigma\ \frac{p(\{\theta\}^{1}_{n_{1}},\Sigma|d)p(d)}{\pi_{\rm PE}(\{\theta\}^{1}_{n_{1}},\Sigma)}p(\Sigma|\Lambda)\prod_{i=1}^{n_{1}}p_{1}(\theta^{1}_{i}|\Lambda)\frac{N_{1}^{n_{1}}}{n_{1}!}e^{-N_{1}}.$ (3.17)
@@
where the sum runs over $N_{s}$ samples of $(\{\theta\}_{n_{1}},n_{1},\Sigma)$ obtained with RJMCMC.
@@ 5 Conclusions
An alternative to the case discussed here would be to introduce a parametrised resolvability function already at the level of the Global Fit. This would let the prior on the parameters of resolved sourced be adjusted to better represent data in terms of a Gaussian component and resolved sources, rather than using a fixed prior or fixed SNR threshold. Determining which regime enables precise astrophysical inference, and developing the tools required for realistic LISA data, will be the subject of future work.
Finally, how our approach can be extended to include other sources, in particular massive black hole binaries and whether injection campaigns as used by the LVK [92] will be needed will also be the subject of future work.
@@ Acknowledgments
We are thankful to G. Faggioli for his contribution in the early stages of this work and to S. Babak, R. Buscicchio and N. Karnesis for fruitful discussions. A.T. is supported by MUR Young Researchers Grant No. SOE2024-0000125, ERC Starting Grant No. 945155–GWmining, Cariplo Foundation Grant No. 2021-0555, MUR PRIN Grant No. 2022-Z9X4XS, Italian-French University (UIF/UFI) Grant No. 2025-C3-386, MUR Grant “Progetto Dipartimenti di Eccellenza 2023-2027” (BiCoQ), and the ICSC National Research Centre funded by NextGenerationEU. Appendix
@@ Appendix A LVK limit
$p(d,\{\theta\}^{1}_{K}|\Lambda),\{\theta\}^{2}_{K}|\Lambda)\propto\prod_{i=1}^{K}\left[f_{i}p_{1}(d_{c,i}|\Lambda)+(1-f_{i})g_{i}p(d_{c,i}|\Lambda)+(1-f_{i})(1-g_{i})p_{n}(d_{c,i})\right],$ (A.6)
@@
$p(d,\{\theta\}^{1}_{K}|\Lambda),\{\theta\}^{2}_{K}|\Lambda)$
$\propto\prod_{i=1}^{K}\left[\frac{N_{1}}{K}p_{1}(d_{c,i}|\Lambda)+\left(1-\frac{N_{1}}{K}\right)\frac{N_{2}}{K}p_{2}(d_{c,i}|\Lambda)+\left(1-\frac{N_{1}}{K}\right)\left(1-\frac{N_{2}}{K}\right)p_{n}(d_{c,i})\right].$ (A.8)

### v3

@@ 1 Introduction
The hierarchical Bayesian formalism [1, 2, 3, 4] is now routinely applied in gravitational-wave (GW) astronomy to deconvolve measurement errors and selection effects in order to measure the global properties of source populations. It is used extensively in the analysis of LIGO/Virgo/KAGRA (LVK) [5, 6, 7, 8] observations to infer the astrophysical properties of stellar-mass compact objects [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22], to measure cosmological parameters [23, 24, 25], and to perform tests of General Relativity [26, 27, 28, 29]. A similar framework is also applied to pulsar-timing-array (PTA) [30, 31, 32] data in order to infer the properties of supermassive black holes in the late Universe [33, 34, 35]. Future detectors, such as the Laser Interferometer Space Antenna (LISA) [36, 37], will require a new population inference paradigm, however. Crucially, while current LVK observations consist exclusively of individually resolved mergers of stellar-mass compact objects, and PTAs observe a stochastic background generated by the superposition of inspiralling supermassive black hole binaries, neither of these regimes will apply to LISA, which will instead observe a superposition of resolved and unresolved signals from the same astrophysical population.
@@
The hierarchical Bayesian formulations for population inference usually treat either (i) a set of statistically independent, individually resolved events, or (ii) a stochastic signal that can be related to the underlying source population. However, they do not accommodate situations where a single population contributes simultaneously to a resolvable population and to an unresolved confusion background - precisely the case for LISA GBs, and potentially for EMRIs. A formalism that includes both resolved events and stochastic backgrounds was proposed in the LVK context [74], but it does not treat resolved and unresolved sources in a fully consistent way. As a consequence, it does not account for the correlations that naturally emerge in a LISA-like Global Fit, particularly the fact that the number of resolvable events is not fixed and that resolving more or fewer sources directly impacts the inferred background. As LISA data-analysis tools are being developed, it is becoming essential to establish a unified formalism capable of treating such populations, both to guide methodological development and to define the data products that should be provided by the collaboration. This issue is also relevant for third-generation ground-based detectors, which will come online in the next decade, such as the Einstein Telescope [75] and Cosmic Explorer [76], which are expected to detect an astrophysical stochastic background [77, 78, 79, 80, 81, 82] in addition to individual events, and for pulsar timing arrays, which are expected to resolve individual sources in the near future [83, 84, 85, 86], particularly once the Square Kilometre Array [87] becomes operational.
In this paper, we develop a more general hierarchical Bayesian framework capable of processing the output of a Global Fit in order to describe populations that generate both a stochastic background and individually resolvable signals. In short, the problem can be thought of as involving two populations: one that contributes to the background and another that gives rise to resolvable sources. Crucially, these populations overlap in parameter space, and we derive how to account for this in a self-consistent manner when evaluating the population likelihood in Sec. 3, after reviewing the approaches that are currently employed for LVK and PTA analyses in Sec. 2. In the appendix, we discuss the conditions under which our formulation reduces to the standard hierarchical Bayesian approach and how it differs from that of [74]. Finally, using a toy model Global Fit, we illustrate how the framework can be applied in practice and highlight the interplay between the stochastic and resolved components in constraining the underlying astrophysical population. We stress that the toy model and its implications for LISA data analysis are very understandable even without the derivations presented in Sec. 3. We therefore encourage readers interested in LISA, but less in the statistical framework for population inference, to read this section independently of Sec. 3.
@@ 2 Current hierarchical Bayesian formalism
In the above, the second term dependence on the full data is only an approximation, as it includes also all of the resolved-source containing segments, $\{d_{c}\}_{n_{\rm det}}$. This is acceptable for current ground-based detectors because data chunks containing detectable events are much rarer than those without. We can therefore assume that these rare events do not affect the reconstruction of the stochastic noise, and conversely, that the noise can be considered known when these events are observed, and does not need to be constrained jointly. However, as we will show in App. A, this expression does not fully incorporate the information of resolved and unresolved sources self-consistently . It is a good approximation for the current situation, where we can only place upper bounds on the background. We also refer to Sec. IV of the Supplemental Material of Ref. [88] for a discussion on the validity of this approximation for third-generation ground-based detectors.
@@ 3 General hierarchical Bayesian formalism
$p(d,\{\theta\}_{n},n,\Sigma|\Lambda)\propto p(d|\{\theta\}_{n},n,\Sigma)p(\Sigma|\{\theta\}_{n},n,\Lambda)\prod_{i=1}^{n}N(\theta_{i}|\Lambda)e^{-N}.$ (3.1)
The above equation states that the likelihood for the data $d$ depends on the events and on the noise properties, and that the noise properties can also depend on the events drawn from $N(\theta|\Lambda)$. We will ultimately marginalise over $\{\theta\}_{n}$, $n$ and $\Sigma$ to obtain the population likelihood. The noise parameters $\Sigma$ can also contain information about quantities that do not depend on the source population, e.g., instrumental noise, but we keep the same notation for generality. For this same reason, the probability of $\Sigma$ is also conditioned on $\Lambda$, allowing us to capture both the components of the noise that depend on the population through $(\{\theta\}_{n},n)$ and those that are independent of the astrophysical model being considered. Note that throughout the derivation, we make no assumption about the likelihood, $p(d|\{\theta\}_{n},n,\Sigma)$, being Gaussian.
@@ 3.1 Separable case
$p(d|\Lambda)\propto$ $\sum_{n_{1}=0}^{+\infty}\sum_{n_{2}=0}^{+\infty}\frac{N^{n_{1}}N^{n_{2}}}{n_{1}!n_{2}!}e^{-N}\int{\rm d}\{\theta\}^{1}_{n_{1}}{\rm d}\{\theta\}^{2}_{n_{2}}{\rm d}\Sigma\ p(d|\{\theta\}^{1}_{n_{1}},n_{1},\Sigma)p(\Sigma|\{\theta\}^{2}_{n_{2}},n_{2},\Lambda)$
$\times\prod_{i=1}^{n_{1}}p(\theta^{1}_{i}|\Lambda)\prod_{j=1}^{n_{2}}p(\theta^{2}_{j}|\Lambda).$ (3.2)
We split the product over $n$ sources as a product over $n_{1}$ resolvable and $n_{2}$ unresolvable sources, and divide by the corresponding factorial terms to avoid overcounting when marginalising. This is because, in our approach, resolvable and unresolvable sources are distinguishable from each other, but sources within each class are indistinguishable once we marginalise. Defining
@@
$p(d|\Lambda)\propto$ $\sum_{n_{1}=0}^{+\infty}\sum_{n_{2}=0}^{+\infty}\frac{N_{1}^{n_{1}}N_{2}^{n_{2}}}{n_{1}!n_{2}!}e^{-N_{1}-N_{2}}\int{\rm d}\{\theta\}^{1}_{n_{1}}{\rm d}\{\theta\}^{2}_{n_{2}}{\rm d}\Sigma\ p(d|\{\theta\}^{1}_{n_{1}},n_{1},\Sigma)p(\Sigma|\{\theta\}^{2}_{n_{2}},n_{2},\Lambda)$
$\times\prod_{i=1}^{n_{1}}p_{1}(\theta^{1}_{i}|\Lambda)\prod_{j=1}^{n_{2}}p_{2}(\theta^{2}_{j}|\Lambda),$ (3.6)
where we divided the $p(\theta^{1,2})$ terms by the normalisation of Eq. (3.3) to obtain $p_{1,2}(\theta^{1,2})$, and absorbed the corresponding normalisation factors into the $N^{n_{1,2}}$ terms.
@@ 3.2 General case
$p(d|\Lambda)=\sum_{n_{1}=0}^{+\infty}\sum_{n_{2}=0}^{+\infty}\frac{N^{n_{1}}N^{n_{2}}}{n_{1}!n_{2}!}e^{-N}\int{\rm d}{\theta}^{1}_{n_{1}}{\rm d}{\theta}^{2}_{n_{2}}{\rm d}\Sigma\ p(d|\{\theta\}^{1}_{n_{1}},n_{1},\Sigma)p(\Sigma|\{\theta\}^{2}_{n_{2}},n_{2},\Lambda)$
$\times\prod_{i=1}^{n_{1}}p(\theta^{1}_{i}|\Lambda)\alpha(\theta^{1}_{i},\Lambda)\prod_{j=1}^{n_{2}}p(\theta^{2}_{j}|\Lambda)(1-\alpha(\theta^{2}_{j},\Lambda)).$ (3.11)
@@
$p(\Sigma|n_{2},\Lambda)=\int{\rm d}\{\theta\}^{2}_{n_{2}}\ p(\Sigma|\{\theta\}^{2}_{n_{2}},n_{2},\Lambda)\prod_{j=1}^{n_{2}}p_{2}(\theta^{2}_{j}|\Lambda),$ (3.14)
@@
$p(d|\Lambda)\propto\sum_{n_{1}=0}^{+\infty}\int{\rm d}\{\theta\}^{1}_{n_{1}}{\rm d}\Sigma\ p(d|\{\theta\}^{1}_{n_{1}},n_{1},\Sigma)p(\Sigma|\Lambda)\prod_{i=1}^{n_{1}}p_{1}(\theta^{1}_{i}|\Lambda)\frac{N_{1}^{n_{1}}}{n_{1}!}e^{-N_{1}}.$ (3.16)
This is main result of this paper, the general equation to be used for hierarchical Bayesian inference for LISA data. We will illustrate this with a single population chosen to represent GBs, but this can be readily extended to multiple populations, for example massive black hole binaries and EMRIs. We stress that our formalism does not rely on the likelihood used in the Global Fit ($p(d|\{\theta\}^{1}_{n_{1}},n_{1},\Sigma)$) being Gaussian, and heavier-tailed likelihoods could be used [90, 91]. In practice, when running a Global Fit with RJMCMC, we have samples from $p(\{\theta\}^{1}_{n_{1}},n_{1},\Sigma|d)$. Using Bayes’ theorem, and denoting by $\pi_{\rm PE}(\{\theta\}^{1}_{n_{1}},n_{1},\Sigma)$ the prior used to obtain the posterior during initial sampling, we have
$p(d|\Lambda)\propto\sum_{n_{1}=0}^{+\infty}\int{\rm d}\{\theta\}^{1}_{n_{1}}{\rm d}\Sigma\ \frac{p(\{\theta\}^{1}_{n_{1}},n_{1},\Sigma|d)p(d)}{\pi_{\rm PE}(\{\theta\}^{1}_{n_{1}},n_{1},\Sigma)}p(\Sigma|\Lambda)\prod_{i=1}^{n_{1}}p_{1}(\theta^{1}_{i}|\Lambda)\frac{N_{1}^{n_{1}}}{n_{1}!}e^{-N_{1}}.$ (3.17)
@@
where the sum runs over $N_{s}$ samples of $(\{\theta\}_{n_{1}},n_{1},\Sigma)$ obtained with RJMCMC. We note that here we are implicitly assuming that the prior on individual event parameters used for the initial parameter estimation was fixed. An alternative strategy would be to do the initial parameter estimation including a population model, i.e., introducing hyperparameters into the individual source priors and simultaneously sampling on both the hyperparameters and the event parameters, as recently suggested in the LVK context [92].
@@ 5 Conclusions
An alternative to the case discussed here would be to introduce a parametrised resolvability function already at the level of the Global Fit. This would let the prior on the parameters of resolved sourced be adjusted to better represent data in terms of a Gaussian component and resolved sources, rather than using a fixed prior or fixed SNR threshold. Beyond this, performing a first population inference in the Global Fit using parametrised priors would also improve the efficiency of the Monte Carlo estimators used to evaluate the population likelihood, see Eq. (3.18), one of the bottlenecks of current hierarchical analyses on LVK data [98, 99]. Samples obtained using a particular parametrised population model can be reweighted to alternative population models in post-processing in various ways. A first approach is to use Eq. (3.18), allowing the prior on individual-event parameters to vary rather than remain fixed, which is not a problem within our formalism. Another possibility is to use a remapping technique. The full Global Fit output could be remapped by assigning pairs of individual-parameter and hyperparameter samples from the initially fitted model to the target model, based on a suitable distance metric between the target and first-fitted models. More simply, one could marginalise over individual event parameters, as usually done in population studies, and remap only the population parameters, following methods recently developed in the LVK context [100, 101] to map flexible non-parametric reconstructions onto parametric or astrophysically motivated models. Determining which regime enables precise astrophysical inference, and developing the tools required for realistic LISA data, will be the subject of future work.
Finally, how our approach can be extended to include other sources, in particular massive black hole binaries and extreme-mass-ratio inspirals (EMRIs) and whether injection campaigns as used by the LVK [102] will be needed will also be the subject of future work. In considering other source types, an additional complication arises due to the size of the signal parameter space. The formalism here relies on the fact that it is possible to obtain converged RJMCMC samples over the full model space. In practice, the EMRI parameter space is so large that this is unlikely to be possible [103, 104]. Instead, the analysis of the data will be separated into a search phase where EMRIs are identified in the data, followed by a parameter inference phase where the posteriors are obtained while simultaneously adjusting the parameters of all sources in the global fit. In such a set up, the separation between resolved sources and background is not purely driven by the data, but also by the effectiveness of the search pipeline. This inference model can also be accommodated in our pipeline, with the resolvability function now determined by the probability that the search identifies a source with given parameters. This can be determined on the fly as the search proceeds, or by an injection campaign, and then kept fixed or allowed to adjust slightly during the refinement phase. While technically there is a deterministic mapping between the observed data and the number of EMRIs that will be found and then fitted, the size of the EMRI parameter space is large enough that it should be possible to treat this probabilistically (as in the framework here), with the implied average over data realisations replaced by an average over the parameter space. A full exploration of the EMRI population fitting problem, and introduction of any necessary refinements, should be the subject of future studies, but the formalism we have described here should provide a valid framework also for that work.
@@ Acknowledgments
We are thankful to G. Faggioli for his contribution in the early stages of this work and to S. Babak, R. Buscicchio, A. Criswell and N. Karnesis for fruitful discussions. A.T. is supported by MUR Young Researchers Grant No. SOE2024-0000125, ERC Starting Grant No. 945155–GWmining, Cariplo Foundation Grant No. 2021-0555, MUR PRIN Grant No. 2022-Z9X4XS, Italian-French University (UIF/UFI) Grant No. 2025-C3-386, MUR Grant “Progetto Dipartimenti di Eccellenza 2023-2027” (BiCoQ), and the ICSC National Research Centre funded by NextGenerationEU. Appendix
@@ Appendix A LVK limit
$p(d|\Lambda)\propto\prod_{i=1}^{K}\left[f_{i}p_{1}(d_{c,i}|\Lambda)+(1-f_{i})g_{i}p_{2}(d_{c,i}|\Lambda)+(1-f_{i})(1-g_{i})p_{n}(d_{c,i})\right],$ (A.6)
@@
$p(d|\Lambda)\propto\prod_{i=1}^{K}\left[\frac{N_{1}}{K}p_{1}(d_{c,i}|\Lambda)+\left(1-\frac{N_{1}}{K}\right)\frac{N_{2}}{K}p_{2}(d_{c,i}|\Lambda)+\left(1-\frac{N_{1}}{K}\right)\left(1-\frac{N_{2}}{K}\right)p_{n}(d_{c,i})\right].$ (A.8)